OpenStudy (anonymous):

how do I change the limits? I thought I was suppose to 2 and 1 for \[\theta\] \[\int_{1}^{2} \frac{\sqrt(x^2 -1)}{x}dx\] \[x = sec\theta\] \[dx = tan\theta \sec \theta d\theta\]

OpenStudy (anonymous):

\[=\int tan^2 \theta d\theta\]

OpenStudy (anonymous):

\[\int \frac{\sqrt(sec^2 \theta -1)}{sec\theta}* tan\theta sec\theta d \theta\] \[\int \frac{tan\theta}{sec\theta}* tan\theta sec\theta d \theta\]

OpenStudy (anonymous):

i forgot how to change the limit....thought I had to plug in 2 and 1 for theta ...but that's wrong

OpenStudy (turingtest):

\[2=\sec\theta=\frac1{\cos\theta}\implies\cos\theta=\frac12\implies\theta=\frac\pi3\]\[1=\sec\theta\implies\cos\theta=1\implies\theta=0\]

OpenStudy (turingtest):

we are just putting the numbers in and finding out what theta is based on our substitution

OpenStudy (anonymous):

life saver! thanks!

OpenStudy (turingtest):

welcome :)

OpenStudy (anonymous):